OpenStudy (anonymous):

how many solutions does the following equations have? y=-1/2x+2 and 3x+6y=12

OpenStudy (compassionate):

\(\Huge\bf Welcome~to~ \color{#00B4ff}{Open}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge\bf\color{#7cc517}{Study~!!}~\) We will need to solve the second equation and put in slope-intercept form. Once we do that, we graph them. \[y = -\frac{ 1 }{ 2 }x + 2\]\[y = -\frac{ 1 }{ 2 }x - 2\] As a rule of thumb: Same slope = No solutions. The lines are parallel and will never intersect. Same equation = Infinite solutions. The lines are the same.  These two lines will never intersect No solution. Thank you for visiting OpenStudy! Please - come back again :) \(\bf\Large\color{#00B4ff}{Open}\color{#7cc517}{Study}\color{}{~Ambassador}\color{#872af0}{~~Ⓐ}\color{#478903}{~Compassionate}\)